Steam enters a turbine stage with an enthalpy of 3628 kJ/kg at 70 m/s and leaves at the

same stage with an enthalpy of 2846 kJ/kg at 124 m/s. Calculate the work per kg done by

the steam.


1
Expert's answer
2021-11-12T10:02:23-0500

"H_1=3628kj/kg\\H_2=2846kj/kg\\v_1=70m/sec\\v_2=124m/sec"

"E_{in}=E_{out}"

"KE_1+H_1=KE_2+H_2+w"

"w=H_1-H_2+\frac{1}{2}(mv_1^2-mv_1^2)"


"w=3628-2846+\frac{1}{2\times1000}(124^2-70^2)"

"w=782+5.238=0.787238kJ\/kg"